OpenStudy (anonymous):

Find the sum of the following infinite geometric series, if it exists: 1/2 + 1/4 + 1/8 + 1/16 + ...

OpenStudy (kropot72):

The formula for the sum to infinity is \[S _{\infty}=\frac{a}{1-r}\] In the question a = 1/2 and r = 1/2, Therefore you just need to plug the values into the formula.